OpenStudy (anonymous):

I'm pretty sure the answer to this is D, but the integral doesn't come out quite right. 6. The function g(x) = 12x2 − sin x is the first derivative of f(x). If f(0) = −2, what is the value of f(2π)? A. 30π + 2 B. 8π2 − 2 C. 4π3 + 8 D. 32π3 − 2 E. 2π2 + 5

OpenStudy (cwrw238):

integral of g(x) = 4x^3 + cos x + c = f(x) f(0) = -2 = cos 0 + c so c = -2-1 = -3 f(2pi) = 32pi^3 + cos(2pi) - 3 = 32pi^3 - 2

OpenStudy (cwrw238):

yup D

OpenStudy (cwrw238):

( as cos(2pi) = 1 )

OpenStudy (anonymous):

ohhh.....that's why. I got the same thing you did for the integral. Thanks!

OpenStudy (cwrw238):

yw